# ARBoids: Adaptive Residual Reinforcement Learning With Boids Model for Cooperative Multi-USV Target Defense

### v2

@@ Abstract
Index Terms:
Marine robotics, reinforcement learning, target defense problem.
@@ I Introduction
Unmanned surface vehicles (USVs), with their ability to autonomously navigate and cooperate in complex environments, have garnered significant attention in the marine robotics community[1, 2, 3]. Among diverse applications, the target defense problem (TDP) emerges as a critical and challenging task, wherein one or more defending USVs are deployed to intercept an attacking USV before it breaches a designated target region. This problem holds practical significance in domains such as maritime surveillance, border security, and the protection of critical marine infrastructure[4, 5, 6].
The TDP is inherently complex due to its dynamic and adversarial nature, requiring defenders to make real-time decisions while anticipating the attacker’s evasive maneuvers. Prior research has made substantial strides in addressing this problem. For instance, Shishika et al. [7] analyzed the TDP under partial information constraints and derived an optimal defensive strategy. Guo et al. [8] designed a neural network-based controller for cooperative capture. Escobedo et al. [9] introduced force-based control with time-varying parameters. Despite these advances, existing approaches largely assume that defenders possess equal or superior agility, thereby limiting their effectiveness in scenarios involving highly maneuverable attackers.
@@
Analytical approaches based on differential game theory and related mathematical frameworks have been extensively explored in recent years [10, 11, 12, 13]. While these methods offer elegant closed-form solutions for idealized adversarial scenarios, they rely on simplifying assumptions such as linearized dynamics and idealized constraints, which do not capture the full complexity of real-world USV systems. In practice, factors such as hydrodynamic interactions, actuator limitations, and environmental disturbances significantly influence USV behavior and are often neglected in theoretical models. Consequently, despite their theoretical merit, these approaches have limited applicability in practical interception scenarios.
Alternatively, force-based methods [14, 15] employ virtual potential fields to guide robot behaviors, achieving cooperative behaviors through local interactions. However, these approaches require extensive parameter tuning and exhibit limited robustness when confronting intelligent or highly maneuverable adversaries. Recently, deep reinforcement learning (DRL) has emerged as a powerful tool for learning complex control policies in high-dimensional, continuous domains [16, 17, 18, 19, 20, 21]. DRL enables agents to acquire diverse behaviors through interaction with simulated environments. Nevertheless, DRL suffers from well-documented limitations — such as low sample efficiency, poor generalization, and unstable training process [22, 23] — that hinder its effectiveness in multi-USV target defense contexts.
To tackle these challenges, this letter proposes ARBoids, a novel adaptive residual reinforcement learning (RL) framework that integrates DRL with the biologically inspired, force-based Boids model [24, 25]. Specifically, the Boids model serves as a computationally efficient baseline policy for coordination, while DRL learns an adaptive residual policy that refines the defenders’ actions. This hybrid approach enables defenders to maintain efficient and scalable behaviors via Boids while dynamically adjusting to complex and challenging scenarios through the learned residual. As a result, ARBoids is particularly effective against attackers with superior maneuverability, combining the robustness of force-based control with the adaptability of learning-based strategies.
@@
  • The proposed ARBoids framework is extensively evaluated in a high-fidelity simulation environment, and its superior performance, scalability, and robustness are demonstrated.
@@ II-A USV Motion Model
In this letter, we employ the widely adopted planar three-degree-of-freedom dynamic model to describe the motion of both the attacking and defending USVs [26], i.e.,
$\bm{M}_{RB}\dot{\bm{\nu}}+\bm{C}_{RB}(\bm{\nu})\bm{\nu}+\bm{M}_{A}\dot{\bm{\nu}}_{r}+\bm{N}(\bm{\nu}_{r})\bm{\nu}_{r}$ (1)
$=\bm{\tau}+\bm{\tau}_{\text{wind}}+\bm{\tau}_{\text{wave}}.$
Here, $\bm{\nu}$ denotes the USV’s velocity in the body-fixed frame, comprising surge, sway, and yaw rate components. $\bm{\nu}_{r}$ represents the relative velocity, accounting for the effects of flow currents. The matrices $\bm{M}_{RB}$, $\bm{C}_{RB}(\bm{\nu})$, $\bm{M}_{A}$, and $\bm{N}(\bm{\nu}_{r})$ represent the rigid-body mass, Coriolis and centripetal, added mass, and hydrodynamic damping terms, respectively. The control input $\bm{\tau}$ corresponds to the forces and moments generated by the USV’s thrusters, while $\bm{\tau}_{\text{wind}}$ and $\bm{\tau}_{\text{wave}}$ denote external disturbances induced by wind and waves.
@@ II-B USV Target Defense Problem
This letter addresses a cooperative target defense problem involving a team of homogeneous USVs tasked with protecting a designated region from a more agile attacking USV, as illustrated in Fig. 1. Given that individual defenders lack the capability to intercept the attacker independently, their objective is to collaboratively capture the attacker through coordinated actions. A trial is considered successful if either of the following conditions is met: (1) the distance between the attacker and at least one defender falls below a predefined capture radius ($d_{i,A}<r_{\text{cap}}$), at any point during the engagement, or (2) the attacker fails to reach the target region within a fixed duration $T_{\text{total}}$. Conversely, a trial is deemed a failure if the attacker breaches the boundary of the circular target region, described by $\|\bm{x}_{A}\|\leq r_{0}$, where $\bm{x}_{A}$ represents the position vector of the attacking USV. Additionally, defenders must maintain safe separation throughout the mission. A trial is also classified as a failure if the distance between any two defenders violates the collision threshold ($d_{i,j}\leq r_{\text{collision}}$).
At the initial stage of each trial, the defending USVs are positioned within the protected region, while the attacking USV approaches from a randomly chosen direction within the defenders’ sensing range. The propulsion of both the defenders and the attacker is subject to bounded thrust constraints, defined as $\tau_{D}\in[\tau^{\text{min}}_{D},\tau^{\text{max}}_{D}]$ and $\tau_{A}\in[\tau^{\text{min}}_{A},\tau^{\text{max}}_{A}]$, respectively. Here, $\tau^{\text{max}}$ represents the maximum forward thrust, while $\tau^{\text{min}}$ denotes the maximum backward thrust. The attacker is granted superior maneuverability, satisfying $\tau^{\text{min}}_{A}<\tau^{\text{min}}_{D}$ and $\tau^{\text{max}}_{A}>\tau^{\text{max}}_{D}$, thereby enabling greater acceleration, faster maximum speed, and increased maximum angular velocity. For simplicity, we assume a constant ratio between the attacker’s and defender’s thrust limits, defined as the attacker’s agility level, i.e.,
@@
The USV target defense problem is then formulated as an optimization problem that aims to determine a sequence of defender actions $\{\bm{\tau}_{D}(t)\}$, which maximizes the probability of task success. This task is subject to the USV dynamics (1), the agility level (2), and additional constraints that govern collision avoidance and thrust limitations. The optimization problem is expressed as
@@
s.t. $d_{i,j}>r_{\text{collision}},\ \forall i\neq j,$
$\tau_{D}\in[\tau^{\text{min}}_{D},\tau^{\text{max}}_{D}],\;\tau_{A}\in[\tau^{\text{min}}_{A},\tau^{\text{max}}_{A}],$
@@ III Design of ARBoids Method
To address the TDP challenge, this letter proposes ARBoids, an hybrid policy that adaptively integrates deep reinforcement learning (DRL) with the biologically inspired, force-based Boids model. Specifically, the Boids model provides a computationally efficient baseline for multi-agent coordination, while the DRL component learns a residual policy that adaptively refines and optimizes the defenders’ actions.
@@ III-A Adaptive Residual Policy
The standard residual policy (RP) framework integrates a baseline policy with a learned residual or correction term, enabling the agent to refine and optimize the baseline behavior[22]. Formally, it is defined as
@@
This structure improves training efficiency and stability by leveraging prior knowledge. However, its performance is tightly coupled with the quality of the baseline policy, as the DRL and baseline components contribute equally regardless of context.
@@
where $\theta(s)\in[0,1]$ is dynamically computed to balance the contributions of $\pi_{\text{DRL}}$ and $\pi_{\text{Boids}}$. Notably, when $\theta(s)=0$, the policy falls back entirely on the Boids model, and when $\theta(s)=1$, it relies solely on the DRL agent. The optimal value of $\theta(s)$ is obtained by maximizing the expected reward for the given state $s$
$\theta^{*}=\arg\max_{\theta\in[0,1]}R\Big{(}s,\theta\cdot\pi_{\text{DRL}}(s)+(1-\theta)\cdot\pi_{\text{Boids}}(s)\Big{)},$ (6)
where $R(\cdot)$ denotes the reward function. This optimization ensures that the ARBoids always selects the more advantageous policy or an effective combination thereof. Thus, the obtained reward satisfies
$R\big{(}s,\pi_{\text{ARBoids}}(s)\big{)}\geq\max\Big{(}R\big{(}s,\pi_{\text{DRL}}(s)\big{)},R\big{(}s,\pi_{\text{Boids}}(s)\big{)}\Big{)}.$ (7)
This guarantees that the ARBoids performs at least as well as the better of the two individual policies for any given state $s$.
@@ III-B Soft Actor Critic
We adopt DRL as the learning baseline and cast the TDP as a Markov Decision Process (MDP), defined by the 5-tuple $(\mathcal{S},\mathcal{A},R,\mathcal{P},\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $R$ the reward function, $\mathcal{P}$ the state-transition probabilities, and $\gamma$ the discount factor. The objective of the DRL agent is to maximize the expected cumulative reward, expressed as $\sum_{t=0}^{T}\gamma^{t}r_{t}$, where $r_{t}$ is the reward obtained at time step $t$. We employ the soft actor-critic (SAC) algorithm [27], a state-of-the-art method known for its robustness and sample efficiency. Unlike conventional DRL algorithms, SAC incorporates an entropy-maximization framework, optimizing the following objective
@@
where $\mathcal{H}(\pi(\cdot|s_{t}))$ denotes the entropy of the policy $\pi$, and $\alpha$ is a temperature parameter balancing exploration and exploitation. This formulation encourages diverse action selection, enhances robustness to dynamic uncertainties, and accelerates policy convergence. Furthermore, our learning framework follows the Centralized Training Decentralized Execution (CTDE) paradigm, combined with parameter sharing. A single shared policy is trained using experiences aggregated from all homogeneous defenders, facilitating efficient data utilization and accelerated convergence. During execution, each agent selects actions autonomously based on its own local observations, yielding a fully decentralized control strategy.
@@ III-C Boids Model
The Boids model, originally proposed by C. W. Reynolds[24], is a biologically inspired computational framework for simulating the collective behavior of swarm-like agents, such as bird flocks or fish schools. This decentralized model governs the behavior of each agent using three fundamental rules — separation, alignment, and cohesion. To enable pursuit of the attacker in our context, we extend the standard Boids model by introducing an additional attraction rule. The mathematical formulations employed in this letter are defined as follows
@@
    where $k_{\text{att}}$ is the attraction weight, and $\bm{x}_{A}$ represents the position of the attacker.
The total virtual force acting on defender $i$ is the weighted sum of all individual components, i.e.,
@@ III-D State Representation
Figure 2: Neural network architecture of the proposed ARBoids method. The network is composed of three main modules: (i) the state embedding module, which processes the state component $s_{i,AT}$ and $s_{i,\text{Boids}}$ independently through fully-connected (Fc) layers and encodes $s_{i,D}$ using a mean observation embedding technique[28]; (ii) the decision module, which utilzes three Fc layers to process the concatenated 640-dimensional feature vector and generate the DRL action $a_{\text{DRL}}$; and (iii) the adapter module, which integrates $a_{\text{DRL}}$, the Boids action $a_{\text{Boids}}$, and the hidden feature from the decision module. The adapter outputs a weighting coefficient $\theta$, which is linearly mapped to the range $[0,1]$, enabling an adaptive balance between $a_{\text{DRL}}$ and $a_{\text{Boids}}$ as formulated in (5).
Figure 3: State space components $s_{i,D}$ and $s_{i,AT}$ for each defender $i$, encompassing information about other defenders ($D$), the attacker ($A$), and the target region ($T$). All observations are expressed in local coordinates.
As shown in Fig. 2, the state space of defender $i$ — assuming a total of $n$ defenders — is composed of three components and is defined as $s_{i}=\left[s_{i,D},s_{i,AT},s_{i,\text{Boids}}\right]$. The first component, $s_{i,D}=\left[\phi_{i,1},d_{i,1},\dots,\phi_{i,n},d_{i,n}\right]$, captures the relative bearing and distance observations of other defenders, as illustrated in Fig. 3. Its dimensionality scales with the number of defenders. The second component, $s_{i,AT}=\left[\phi_{i,A},d_{i,A},\phi_{i,T},d_{i,T}\right]$, encodes the relative positions of the attacker and the target. To better approximate real-world sensing conditions, we incorporate observation noise into this component to simulate sensor inaccuracies. The third component, $s_{i,\text{Boids}}=\left[\bm{F}_{i,\text{sep}},\bm{F}_{i,\text{ali}},\bm{F}_{i,\text{coh}},a_{i,\text{Boids}}\right]$, contains the virtual forces defined in (13) and the resulting action $a_{i,\text{Boids}}$ generated by the Boids model. This component serves as the baseline policy and implicitly encodes information regarding teammate positions and group formation, thereby assisting the DRL agent in learning a more effective residual policy.

### v3

@@ I Introduction
Unmanned surface vehicles (USVs), with their ability to autonomously navigate and cooperate in complex environments, have garnered significant attention in the marine robotics community[qiao2023survey, lin2025distributional]. Among diverse applications, the target defense problem (TDP) emerges as a critical and challenging task, wherein one or more defending USVs are deployed to intercept an attacking USV before it breaches a designated target region. This problem holds practical significance in domains such as maritime surveillance, border security, and the protection of marine infrastructure[chen2024distributed, fu2024combined, fu2020guarding].
The TDP is inherently complex due to its dynamic and adversarial nature, requiring defenders to make real-time decisions while anticipating the attacker’s evasive maneuvers. Prior research has made substantial strides in addressing this problem. For instance, Shishika et al. [shishika2021partial] analyzed the TDP under partial information constraints and derived an optimal defensive strategy. Guo et al. [guo2024pursuit] designed a neural network-based controller for cooperative capture. Despite these advances, existing approaches largely assume that defenders possess equal or superior agility, thereby limiting their effectiveness in scenarios involving highly maneuverable attackers.
@@
Analytical approaches based on differential game theory and related mathematical frameworks have been extensively explored in recent years[yan2024multiplayer, garcia2020optimal, lee2021guarding]. While these methods offer elegant closed-form solutions for idealized adversarial scenarios, they rely on simplifying assumptions such as linearized dynamics and idealized constraints, which do not capture the full complexity of real-world USV systems. Alternatively, force-based methods[force1, force2, force3], which employ virtual potential fields to guide robotic movement, has achieved cooperative behaviors. However, they exhibit limited robustness when confronting highly maneuverable adversaries.
Recently, deep reinforcement learning (DRL) has emerged as a powerful tool for learning complex control policies in multi-agent systems[lowe2017multi, rashid2020monotonic]. These algorithms usually follow the Centralized Training Decentralized Execution (CTDE) paradigm, which enables agents to leverage global information during training to develop better coordination strategies while maintaining scalability and robustness through decentralized execution. Therefore, multi-agent RL has gained significant popularity in various applications[rl1, rl2, rl3, rl5]. However, these DRL-based methods suffer from well-documented limitations — such as low sample efficiency, poor generalization, and unstable training process — that hinder its effectiveness in multi-USV target defense scenarios. To address these challenges, residual RL has been proposed as hybrid approach that integrates domain knowledge or heuristic controllers with learned policies[johannink2019residual, rana2023residual]. By leveraging a baseline controller to manage well-understood aspects of a task, residual RL allows the learned policy to focus on refining actions for unmodeled or complex dynamics. This combination enhances both sample efficiency and robustness, making residual RL a promising solution for the multi-USV TDP problem.
This letter proposes ARBoids, a novel adaptive residual RL framework that integrates DRL with the biologically inspired, force-based Boids model [reynolds1987flocks, bajec2007computational]. Specifically, the Boids model serves as a computationally efficient baseline policy for coordination, while DRL learns an adaptive residual policy that refines the defenders’ actions. This hybrid approach enables defenders to maintain efficient and scalable behaviors via Boids while dynamically adjusting to complex and challenging scenarios through the learned residual. As a result, ARBoids is particularly effective against attackers with superior maneuverability, combining the robustness of force-based control with the adaptability of learning-based strategies.
@@
  • The proposed ARBoids framework is extensively evaluated in a high-fidelity simulation environment, and its superior performance, generalizability, and robustness are demonstrated.
@@ II-A USV Motion Model
In this letter, we employ Virtual RobotX (VRX)[bingham2019toward], a high-fidelity, Gazebo-based marine simulation platform, as the evaluation environment. The WAM-V USV model provided by VRX serves as the deployment robot. Its motion dynamics are based on the model described in[fossen2011handbook], expressed as
$\bm{M}_{RB}\dot{\bm{\nu}}+\bm{C}_{RB}(\bm{\nu})\bm{\nu}+\bm{M}_{A}\dot{\bm{\nu}}_{r}+\bm{C}_{A}(\bm{\nu}_{r})\bm{\nu}_{r}$ (1)
$+\bm{D}(\bm{\nu}_{r})\bm{\nu}_{r}+\bm{g}(\bm{\eta)}=\bm{\tau}+\bm{\tau}_{\text{wind}}+\bm{\tau}_{\text{wave}}.$
Here, $\bm{\nu}$ and $\bm{\eta}$ are the 6-dimensional velocity and position vectors, while $\bm{\nu}_{r}$ is the vessel velocity relative to the fluid. $\bm{\tau}$ corresponds to the forces and moments generated by USV’s left and right propellers. To simplify control, the propeller angles are fixed at zero, ensuring that the thrust vectors always align with the longitudinal axis of the hull. Yaw motion is achieved by modulating the thrust differential between two propellers. The maximum forward thrust of a single propeller is denoted as $\tau^{\text{max}}$, while the maximum reverse thrust is represented as $\tau^{\text{min}}$. Consequently, the control input for the USV is defined as the thrust values applied to the two propellers. In this study, external forces such as wind and wave disturbances ($\bm{\tau}_{\text{wind}}$ and $\bm{\tau}_{\text{wave}}$) are ignored.
@@ II-B USV Target Defense Problem
The multi-USV target defense involving a team of homogeneous USVs tasked with protecting a designated region from a more agile attacking USV, as illustrated in Fig. 1. We adopt a target-centered coordinate frame in which the target is fixed at the origin, i.e., $\bm{x}_{T}=(0,0)$. A trial is considered successful if either of the following conditions is met: (1) the distance between the attacker and at least one defender falls below a predefined capture radius ($d_{i,A}<r_{\text{cap}}$), at any point during the engagement, or (2) the attacker fails to reach the target region within a fixed duration $T_{\text{total}}$. Conversely, a trial is deemed a failure if the attacker breaches the boundary of the circular target region, described by $\|\bm{x}_{A}\|\leq r_{0}$, where $\bm{x}_{A}$ represents the attacker’s position relative to the target. Additionally, defenders must maintain safe separation throughout the mission. A trial is also classified as a failure if the distance between any two defenders violates the collision threshold ($d_{i,j}\leq r_{\text{collision}}$).
At the initial stage of each trial, the defending USVs are positioned within the protected region, while the attacking USV approaches from a randomly chosen direction within the defenders’ sensing range. The propulsion of both the defenders and the attacker is subject to bounded thrust constraints, defined as $\tau_{D}\in[\tau^{\text{min}}_{D},\tau^{\text{max}}_{D}]$ and $\tau_{A}\in[\tau^{\text{min}}_{A},\tau^{\text{max}}_{A}]$, respectively. The attacker is granted superior maneuverability, satisfying $\tau^{\text{min}}_{A}<\tau^{\text{min}}_{D}$ and $\tau^{\text{max}}_{A}>\tau^{\text{max}}_{D}$, thereby enabling greater acceleration, faster maximum speed, and increased maximum angular velocity. For simplicity, we assume a constant ratio between the attacker’s and defender’s thrust limits, defined as the attacker’s agility level, i.e.,
@@
The multi-USV TDP is formulated as an optimization problem that aims to find a sequence of defender actions $\{\bm{\tau}_{D}(t)\}$, which maximizes the probability of task success. This task is subject to the USV dynamics (1), the agility level (2), and additional constraints that govern collision avoidance and thrust limitations. Furthermore, the optimization problem accounts for a fixed adversarial attacker’s policy $\pi_{A}$, which governs the attacker’s actions based on its observation $\bm{o}_{A}$. The resulting optimization problem is given by
@@
s.t. $\tau_{D}\in[\tau^{\text{min}}_{D},\tau^{\text{max}}_{D}],\quad\tau_{A}\in[\tau^{\text{min}}_{A},\tau^{\text{max}}_{A}],$
$\bm{\tau}_{A}\sim\pi_{A}(\bm{o}_{A}),\quad d_{i,j}>r_{\text{collision}},\ \forall i\neq j,$
@@ III Design of ARBoids Method
We begin by formulating this problem as a Markov Decision Process (MDP), defined by the 5-tuple $(\mathcal{S},\mathcal{A},R,\mathcal{P},\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ is the action space, $R$ is the reward function, $\mathcal{P}$ is the state-transition probabilities, and $\gamma$ is the discount factor.
@@ III-A Adaptive Residual Policy
The standard residual policy (RP) framework integrates a baseline policy with a learned residual or correction term, enabling the agent to refine and optimize the baseline behavior[johannink2019residual]. Formally, it is defined as
@@
This structure effectively leverages the prior knowledge of baseline policy. However, its performance is tightly coupled with the quality of the baseline policy, as the DRL and baseline components contribute equally regardless of context.
@@
where $\theta(s)\in[0,1]$ is dynamically computed to balance the contributions of $\pi_{\text{DRL}}$ and $\pi_{\text{Boids}}$. Notably, when $\theta(s)=0$, the policy falls back entirely on the Boids model, and when $\theta(s)=1$, it relies solely on the DRL policy. We train $\theta(s)$ end-to-end within our DRL framework to maximize expected discounted return. $\theta^{*}(s)$ represents the optimal state-dependent parameter, defined as
$\theta^{*}(s)=\arg\max_{\theta\in[0,1]}Q^{\pi_{\text{ARBoids}}}\Big(s,\pi_{\text{ARBoids}}(s)\Big),$ (6)
where $Q^{\pi}(s,a)$ represents the expected discounted return starting from the state-action pair $(s,a)$ following policy $\pi$, i.e.,
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}\Big|s_{0}=s,a_{0}=a\right].$ (7)
Here, $r_{t+1}$ denotes the reward at time step $t+1$. In practice, $\theta(s)$ is generated by an adapter network, which is trained jointly with the actor-critic architecture. This end-to-end optimization enables the framework to focus on maximizing cumulative rewards, thereby improving long-term performance.
@@ III-B Soft Actor Critic
We employ the soft actor-critic (SAC) algorithm [SAC], a state-of-the-art method known for its robustness and sample efficiency. SAC incorporates an entropy-maximization framework, optimizing the following objective
@@
where $\mathcal{H}(\pi(\cdot|s_{t}))$ denotes the entropy of the policy $\pi$, and $\alpha$ is a temperature parameter balancing exploration and exploitation. This formulation encourages diverse action selection, enhances robustness to dynamic uncertainties, and accelerates policy convergence. Furthermore, our learning framework follows the CTDE paradigm, combined with parameter sharing. A single shared policy is trained using experiences aggregated from all homogeneous defenders, facilitating efficient data utilization and accelerated convergence. During execution, each agent selects actions autonomously based on its own local observations, yielding a fully decentralized control strategy.
@@ III-C Boids Model
The Boids model, originally proposed by C. W. Reynolds[reynolds1987flocks], is a biologically inspired computational framework for simulating the collective behavior of swarm-like agents, such as bird flocks or fish schools. This decentralized model governs the behavior of each agent using three fundamental rules — separation, alignment, and cohesion. To enable pursuit of the attacker in our context, we extend the standard Boids model by introducing an additional attraction rule. The mathematical formulations employed in this letter are defined as follows
@@
    where $k_{\text{att}}$ is the attraction weight.
The total virtual force acting on defender $i$ is the sum of all individual components, i.e.,
@@
The virtual force $\bm{F}_{i}$ is subsequently transformed into the vehicle’s local coordinate frame, and then linearly mapped to generate the thrust commands, which collectively define the Boids action $a_{i,Boids}$. The Boids hyperparameters are initialized based on established guidelines from the literature and fine-tuned in our simulator to provide a stable baseline policy.
@@ III-D State Representation
Figure 2: Architecture of the policy and critic networks for a single defender agent. The policy network comprises three modules: (i) State embedding, which processes the state component $s_{i,AT}$ and $s_{i,\text{Boids}}$ independently via fully-connected (Fc) layers, and encodes $s_{i,D}$ using mean observation embedding[huttenrauch2019deep]; (ii) Decision, which applies three Fc layers to the concatenated feature vector to produce the DRL action $a_{\text{DRL}}$; and (iii) Adapter, which integrates $a_{\text{DRL}}$, the Boids action $a_{\text{Boids}}$, and the decision-module hidden features. The adapter outputs a weighting coefficient $\theta$, linearly mapped to $[0,1]$, to adaptively weight $a_{\text{DRL}}$ and $a_{\text{Boids}}$. The critic network uses modules (i) and (ii), and includes an additional action-embedding layer to encode $a_{\text{DRL}}$ and $\theta$.
Figure 3: (a) State space components $s_{i,D}$ and $s_{i,AT}$ for each defender $i$, encompassing information about other defenders ($D$), the attacker (AA), and the target region ($T$). All observations are expressed in local coordinates. (b) Diagram of components included in the formation reward $r_{\text{form}}$. Here, $\vec{d}_{iA}$ denotes the unit vector from defender $i$ to the attacker, $\vec{d}_{TA}$ is the unit vector from the target to the attacker, and $\vec{d}_{DA}=\sum_{i=1}^{n}\vec{d}_{iA}$ is the cumulative direction vector of all defenders.
As shown in Fig. 2, the state space of defender $i$ — assuming a total of $n$ defenders — is composed of three components and is defined as $s_{i}=\left[s_{i,D},s_{i,AT},s_{i,\text{Boids}}\right]$. The first component, $s_{i,D}=\left[\phi_{i,1},d_{i,1},\dots,\phi_{i,n},d_{i,n}\right]$, captures the relative bearing and distance observations of other defenders, as illustrated in Fig. 3(a). Its dimensionality scales with the number of defenders. The second component, $s_{i,AT}=\left[\phi_{i,A},d_{i,A},\phi_{i,T},d_{i,T}\right]$, encodes the relative positions of the attacker and the target. To better approximate real-world sensing conditions, we incorporate observation noise into this component during training. The third component, $s_{i,\text{Boids}}=\left[\bm{F}_{i,\text{sep}},\bm{F}_{i,\text{ali}},\bm{F}_{i,\text{coh}},a_{i,\text{Boids}}\right]$, contains the virtual forces defined in (13) and the resulting action $a_{i,\text{Boids}}$ generated by the Boids model. This component serves as the baseline policy and implicitly encodes information regarding teammate positions and group formation, thereby assisting the DRL agent in learning a more effective residual policy.